OpenStudy (erinweeks):

What is the energy equivalent of 5.0 kg of mass? Show your work.

OpenStudy (erinweeks):

@e.cociuba @kylees

OpenStudy (e.cociuba):

Is there a problem that goes with this?

OpenStudy (erinweeks):

no this is all it asks

OpenStudy (e.cociuba):

E = mc² = 5 x 9e16 = 45e16 joules I think?

OpenStudy (erinweeks):

no its a midterm can you find someone to help

OpenStudy (e.cociuba):

@ganeshie8 @Meepi @InsanelyChaotic @Laura<3 @Abarnett @Gabylovesyou @antoni7

OpenStudy (erinweeks):

its alright

OpenStudy (anonymous):

E = mc^2, where E is energy in Joules, m is mass in kg, and c is speed of light, 3.0 x 10^8 m/s. E = mc^2 = 5.0kg x (3.0 x 10^8 m/s)^2 = 4.5 x 10^17 kg•m^2/s^2 = 4.5 x 10^17 Joules

OpenStudy (erinweeks):

thank you